OpenStudy (anonymous):

Find the domain restriction. Fill in the table do preserved pairs, and graph the square root function. F(x)=squrt x-3 Table X & Y 3 4 7 12

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf ƒ }\)\(\LARGE\color{blue}{ \bf (x)=\sqrt{x-3} }\) can you complete your table , at first, by simply plugging in the numbers you are given for domain ?

OpenStudy (solomonzelman):

The domain can be all real numbers besides those that are less than 3, do you agree? Becuase if x<3 then f(x)="imaginary number " .

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf f(x)=\sqrt{x-3} }\) if x=3, \(\LARGE\color{blue}{ \bf f(x)=\sqrt{3-3}=\sqrt{0} =0 }\) if x=4, \(\LARGE\color{blue}{ \bf f(x)=\sqrt{4-3}=\sqrt{1} =±1 }\) if x=7, \(\LARGE\color{blue}{ \bf f(x)=\sqrt{7-3}=\sqrt{4} =±2 }\) if x=12, \(\LARGE\color{blue}{ \bf f(x)=\sqrt{12-3}=\sqrt{9} =±3 }\) This are the ranges ( found by plugging in the domains into the function)